Please help me solve this word problem: The sum of the ages of Mark and his father is 58 years. The difference in their ages is 32 years. How old is each person? I tried solving it by doing 2x+32=58-2x. But then I got a fraction for an answer. Thank you for helping me!
Found 2 solutions by Earlsdon, chelcilovesmj1:
Answer by Earlsdon(6287)

Let F = father's age and M = mikes's age.
F+M = 58 and
F-M= 32 Rewrite as:
F = M+32 and substitute into first equation.
M+32+M = 58 Simplify.
2M+32 = 58
2M = 26
M = 13 and
F = 13+32
F = 45

Answer by chelcilovesmj1(3)

Let A = age of dad
B = age of son
It is given that...
A + B = 58
A - B = 32
now add the two equations as a system...and the B cancels out
we are left with
2A = 90
Therefore A = 45
now use one of first equations to find B
45 + B = 58 or 45 - B = 32
either or we find that B = 13
hope this helps :)